OpenStudy (anonymous):

A pack of 52 cards is distributed among 4 players. The one to receive the set with lowest total wins (A=1, J=11, Q=12, K=13). What is the lowest total with which one can win?

OpenStudy (anonymous):

4 aces, twos, threes and a 4 = 4+8+12+4=28

OpenStudy (anonymous):

as 52/4=13 so each player has 13 cards

OpenStudy (anonymous):

thanks